OpenStudy (anonymous):

(Help with extraneous solutions) Find the solution of sqrX+3+4=6 and determine if it is an extraneous solution. x = 1; not extraneous x = 1; extraneous x = 29; extraneous x = 29; not extraneous

OpenStudy (anonymous):

Find the solution of 5sqrx+7= -10 and determine if it is an extraneous solution. x = –3; extraneous x = –3; not extraneous x = 11; extraneous x = 11; not extraneous

OpenStudy (anonymous):

I'm not completely sure that answers my question.

OpenStudy (anonymous):

I think that helps a bit I'm still confused about how to go about solving it sorry

OpenStudy (rainbow_dashie):

oh ok

OpenStudy (rainbow_dashie):

The instructions are precise

OpenStudy (anonymous):

No I get that but I mean I don't quite understand the extraneous part.

OpenStudy (rainbow_dashie):

Oh ok

OpenStudy (anonymous):

Do you know whether it's extraneous or not because I figured out that the answer is 29.

OpenStudy (anonymous):

hey @Angelwings22 , assuming the question is: \[\sqrt{x+3+4}=6\] then: \[\sqrt{x+7}=6\] square both size to get: x+7=36 x=29 now, extraneous solution is a solution that DOES NOT SATISFY the original equation. so we have to plug 29 into the original equation: \[\sqrt{29+3+4}=6\] \[\sqrt{36}=6\] indeed. so this is not extraneous. can you do your second question by yourself ?